At what speed do a bicycle and its rider, with a combined mass of 100 kg, have the same momentum as a 1500kg car traveling at 4.0m/s ?
Answer #1

combined mass = 100 kg

momentum of 1500 kg car = 1500 x 4 = 6000 kgm/sec

6000 = 100 x speed

speed = 60 m/sec
Answer #2

100*V = 1500*4

v= 60 m/s
Answer #3

let the speed be 'v'

=> 1500*4 = 100*v

so v = 60m/s